OpenStudy (anonymous):

please help me (: If the circumference of a circle is 24, and the angle measure of an arc is 120o, which is the length of the arc?

OpenStudy (anonymous):

120=360/3 C=24 Measure of arc=C/3=24/3=8

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

yw